OpenStudy (anonymous):

5 / 4+√5 1) 20-√5 / 11 2) 20-5√5 / 11 3) 5√5 / 4√5+5 4) 5√5 / 9

OpenStudy (anonymous):

multiply 5 / (4+√5) by (4-√5) to get started

OpenStudy (anonymous):

Multiply both nominator and denominator I mean.

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

\[\frac{5(4-\sqrt{5})}{(4+\sqrt{5})(4-\sqrt(5))}=\frac{20-5\sqrt{5}}{16-25}\] Are you still with me?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Well at this point you can pretty easily see that 1,2 and 4 are not correct. so 3 is the answer. It's still not very pretty though.

OpenStudy (anonymous):

Yeah to me its really confusing but i appreciate the help

OpenStudy (anonymous):

Yeah, well, I was going for this representation: \[-\frac{20}{9}+\frac{5\sqrt{5}}{9}\] But that wasn't an option. This way you don't have roots in your denominator.